### v1

@@ 1.1 Key Element Configurations
The reference configuration of an element is that in which the element experiences no strain and no stress. Typically, this is the undeformed configuration of the element, as produced by a mesher, e.g., Gmsh GeuzaineRemacle2009Gmsh or TetGen Si2015TetGen . Associated with this reference configuration there is a reference frame $Ouvw$, whose position and orientation are fixed in space. A point $P$ in the reference configuration of coordinates $(u,v,w)$ is mapped at time $t$ into a different point, which has coordinates $\mathbf{r}(u,v,w;t)=[x(u,v,w;t),y(u,v,w;t),z(u,v,w;t)]^{T}$ in a global reference frame $Oxyz$. The deformation map $\bm{\varphi}$ is the mapping that takes a point in the reference configuration to a point in the current configuration. The deformation map is a function of the time $t$ and the material coordinates $\mathbf{u}$, i.e., $\bm{\varphi}=\bm{\varphi}(\mathbf{u};t)$, where $\mathbf{u}=[u,v,w]^{T}$. The initial configuration is the configuration of the element at time $t=0$. Finally, the parent (or canonical) configuration is used to carry out the numerical integration over the volume of the element. The elements can be iso-parametric, or, for simple cases, the mapping between the parent and the reference configuration can be an affine transformation.
@@ 1.3.2 ANCF 3443 Shell Element
Let $\mathbf{X}_{a}$ and $\mathbf{x}_{a}$ denote the nodal positions in the reference and current configurations, respectively, and let $\mathbf{d}_{a}=\mathbf{x}_{a}-\mathbf{X}_{a}$ be the displacement of node $a$. The isoparametric mappings for the reference position and displacement fields inside the element are
@@
The current position is $\mathbf{x}=\mathbf{X}+\mathbf{u}$, and the deformation gradient in the Total Lagrangian setting is obtained from
$\mathbf{F}(\xi,\eta,\zeta)=\frac{\partial\mathbf{x}}{\partial\mathbf{X}}=\frac{\partial\mathbf{x}}{\partial(\xi,\eta,\zeta)}\left(\frac{\partial\mathbf{X}}{\partial(\xi,\eta,\zeta)}\right)^{-1}.$ (10)
@@ 2 Augmented Lagrangian Formulation
$\quad+h\,\mathbf{J}(\mathbf{q}_{n}+h\mathbf{v})^{T}\bigl(\bm{\lambda}+\rho\,\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})\bigr)\;,$
where $\mathbf{M}$ is a preassembled and constant mass matrix of the dynamic system, $\rho$ is the ALM penalty parameter, and $\mathbf{c}(\mathbf{q})=0$ denotes bilateral constraints with Jacobian $\mathbf{J}=\partial\mathbf{c}/\partial\mathbf{q}$. Note that $\mathbf{c}(\mathbf{q})\in\mathbb{R}^{m}$ is a constraint-space quantity and is mapped to the generalized space through $\mathbf{J}(\cdot)^{T}$ in Eq. (17). We evaluate $\mathbf{J}^{T}(\cdot)$ via a sparse representation of $\mathbf{J}^{T}$, and enforce constraints with an augmented-Lagrangian outer loop that alternates an inner solve for $\mathbf{v}$ with multiplier updates until $\|\mathbf{c}\|_{2}$ meets a prescribed tolerance.
@@
Using $\nabla_{\mathbf{v}}\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})=h\,\mathbf{J}(\mathbf{q}_{n}+h\mathbf{v})$ and the chain rule, one verifies that $\nabla_{\mathbf{v}}\Phi_{\rho}(\mathbf{v},\bm{\lambda})=\mathbf{g}(\mathbf{v},\bm{\lambda})$. Consequently, for fixed multipliers $\bm{\lambda}$, the inner solve $\mathbf{g}(\mathbf{v},\bm{\lambda})=\mathbf{0}$ is equivalent to enforcing stationarity of the augmented cost (18) with respect to $\mathbf{v}$, while the outer ALM loop updates $\bm{\lambda}$ to reduce the constraint violation $\|\mathbf{c}\|_{2}$.
@@ 5.2.1 Kinematics and Strain-Rate Measure in the Reference Configuration
In the Total Lagrangian setting, all kinematic quantities are expressed with respect to the reference configuration. Let $\mathbf{X}$ denote a material point in the reference configuration and $\mathbf{x}(\mathbf{X},t)$ its current position at time $t$. The deformation gradient is
$\mathbf{F}(\mathbf{X},t)=\frac{\partial\mathbf{x}}{\partial\mathbf{X}},\qquad J=\det\mathbf{F}.$ (55)
@@
$\dot{\mathbf{F}}=\frac{\partial\dot{\mathbf{x}}}{\partial\mathbf{X}},$ (57)
@@ 7 Kinematic Constraint Handling in TL-FEA Formulation
Following the taxonomy proposed in jayAllieDan2022rA , we introduce the following four basic geometric constraints: dot-product 1 (DP1), dot-product 2 (DP2), distance (DIST), and pin (PIN). The DP1 constraint brings into play four points: points P and Q that belong to elements $E$ and $F$ on body b, respectively, and points R and T that belong to body c’s elements $G$ and $H$, respectively. This constraint requires that the dot product between the vector from P to Q and the vector from R to T is equal to a user-prescribed function of time, $f(t)$, where $f:\mathbb{R}_{+}\to\mathbb{R}$. In many cases, this function is constant and zero for all time $t$, in which case the constraint becomes a perpendicularity requirement. Using the TL-FEA framework proposed, DP1 assumes the following form:
$\left[{\mathbf{N}}^{b}_{F}(t)\>\mathbf{s}^{b}_{F}(\text{{Q}})-{\mathbf{N}}^{b}_{E}(t)\>\mathbf{s}^{b}_{E}(\text{{P}})\right]^{T}\left[{\mathbf{N}}^{c}_{H}(t)\>\mathbf{s}^{c}_{H}(\text{{T}})-{\mathbf{N}}^{c}_{G}(t)\>\mathbf{s}^{c}_{G}(\text{{R}})\right]-f(t)=0\;.$ (84)
Similar expressions can be written for the other three basic geometric constraints. DP2 is similar to DP1, but it places the point R on the body that points P and Q belong to. The DIST constraint states that the distance between two points (which can belong to the same element, or same body, or different bodies) should assume a strictly positive, user-prescribed value, possibly time-varying. The PIN constraint states that two points (which can belong to the same element, or same body, or different bodies) coincide at all times. Based on this taxonomy, one can build typical engineering joints such as spherical (one PIN constraint), universal (one PIN & one DP1), cylindrical (two DP1 and two DP2), revolute (one PIN and two DP1), translational (three DP1 and two DP2), etc.

### v2

@@ Abstract
1 Background and Related Work
Large-deformation deformable-body dynamics can be posed naturally within a continuum mechanics setting by describing kinematics through the deformation map $\mathbf{r}(\mathbf{X},t)$ and its associated deformation gradient $\mathbf{F}(\mathbf{X},t)=\partial\mathbf{r}/\partial\mathbf{X}$. In a Total Lagrangian (TL) formulation, all strain measures, stress measures, and constitutive responses are expressed with respect to the reference configuration, leading to a conceptually clean separation between geometry (reference domain), kinematics (through $\mathbf{F}$ and its invariants), and material response (through a strain energy density and/or rate-dependent extensions). This deformation-gradient-centric viewpoint is standard in nonlinear finite elements for solids and structures and provides a universal interface for accommodating hyperelasticity, finite-strain viscous regularization, and multi-field extensions while retaining a consistent linearization and tangent structure Bonet2008nonlinear .
Here, $\mathbf{X}$ denotes generic reference/material coordinates. In element-level derivations, we parameterize the reference domain with local coordinates $\mathbf{u}=[u,v,w]^{T}$, so that locally $\mathbf{X}\equiv\mathbf{u}$.
Within flexible multibody dynamics, the Absolute Nodal Coordinate Formulation (ANCF) introduced by Shabana was a key step toward combining large rigid-body motion with large deformation in a non-incremental finite-element setting Shabana1996 . The defining feature of ANCF is the choice of nodal coordinates: instead of interpolating rotations, ANCF employs absolute position vectors and their spatial gradients as nodal degrees of freedom. This coordinate selection yields a polynomial representation of the position field that remains valid under arbitrarily large rotations and deformations, and it enables a constant (configuration-independent) consistent mass matrix for many element families Shabana1998 . In Shabana’s original presentation and in the standard ANCF literature that followed, the position field is written in the familiar FE form $\mathbf{r}(\mathbf{X},t)=\mathbf{S}(\mathbf{X})\mathbf{e}(t)$, where $\mathbf{S}(\mathbf{X})$ is a block-structured shape-function matrix and $\mathbf{e}(t)$ stacks nodal positions and slopes into a generalized coordinate vector shabanaYakoub2001Part1 ; Shabana2023Overview ; RecueroNegrut2016ANCF .
A large body of work has since expanded ANCF element technology and clarified its modeling tradeoffs. Early developments addressed the construction of beam and cable elements that admit large axial and bending deformation while capturing geometrically nonlinear coupling effects without resorting to rotation interpolation berzeri2000 . Subsequent research introduced shear-deformable and higher-order elements and investigated numerical pathologies such as shear and membrane locking, together with formulation-level remedies that retain ANCF kinematics while improving element accuracy and robustness GarciaVallejo2007LockingFree . Comprehensive surveys document these developments and highlight a recurring theme: although ANCF elements come in many variants (beams, plates/shells, solid-like continua), their kinematics can be consistently viewed as a continuum mapping whose primary computational artifact is the deformation gradient and its derived strain measures Gerstmayr2013 ; Otsuka2022RecentAdvances ; Shabana2023Overview .
The present work is built on this lineage but reorganizes the kinematic interpolation to make the deformation-gradient structure explicit. Rather than embedding the spatial basis in a shape-function matrix $\mathbf{S}(\mathbf{X})$ acting on a generalized-coordinate vector, we express the position field as $\mathbf{r}(\mathbf{X},t)=\mathbf{N}(t)\mathbf{s}(\mathbf{X})$, where $\mathbf{s}(\mathbf{X})$ collects scalar basis functions and $\mathbf{N}(t)$ stores the nodal unknowns in a matrix form. This representation is algebraically equivalent to the classical ANCF interpolation (it corresponds to a different arrangement of the same degrees of freedom), yet it changes where the shape functions “live” and, crucially, yields the factorization $\mathbf{F}(\mathbf{X},t)=\mathbf{N}(t)\mathbf{H}(\mathbf{X})$ with $\mathbf{H}(\mathbf{X})=\partial\mathbf{s}/\partial\mathbf{X}$. As a result, element kinematics, stress measures, and internal force contributions can be written uniformly in terms of $\mathbf{F}$ (and, when needed, $\dot{\mathbf{F}}$), which aligns the ANCF description with standard TL continuum mechanics notation and facilitates a unified derivation across element topologies and material models Bonet2008nonlinear ; Shabana2023Overview .
From the standpoint of formulation comparison, the deformation-gradient emphasis is consistent with the direction of modern ANCF literature: reviews stress that many apparent differences between ANCF element variants ultimately reduce to (i) the choice of interpolation spaces for $\mathbf{r}(\mathbf{X},t)$ and its gradients, (ii) the induced strain measures and their linearizations, and (iii) the treatment of cross-section deformation assumptions (e.g., Euler–Bernoulli, Timoshenko-like, or fully continuum) Gerstmayr2013 ; Otsuka2022RecentAdvances . By casting all element expressions through $\mathbf{F}$ in a TL setting, the present framework is positioned as a universal, element-agnostic formulation layer: beams, shells, and solid elements differ primarily through their basis functions and reference-domain parameterizations, while the constitutive and variational machinery remains shared.
@@ 2 Formulation for Deformable Elements
Throughout, $\mathbf{X}$ denotes generic reference/material coordinates. Within an element, we parameterize the reference domain by local coordinates $\mathbf{u}=[u,v,w]^{T}$, so that locally $\mathbf{X}\equiv\mathbf{u}$.
@@ 2.1 Key Element Configurations
The reference configuration of an element is that in which the element experiences no strain and no stress. Typically, this is the undeformed configuration of the element, as produced by a mesher, e.g., Gmsh GeuzaineRemacle2009Gmsh or TetGen Si2015TetGen . Associated with this reference configuration there is a reference frame $Ouvw$, whose position and orientation are fixed in space. A point $P$ in the reference configuration of coordinates $(u,v,w)$ is mapped at time $t$ into a different point, which has coordinates $\mathbf{r}(u,v,w;t)=[x(u,v,w;t),y(u,v,w;t),z(u,v,w;t)]^{T}$ in a global reference frame $Oxyz$. In this work, the deformation map is represented by the position field $\mathbf{r}(\mathbf{u};t)$, where $\mathbf{u}=[u,v,w]^{T}$. The initial configuration is the configuration of the element at time $t=0$. Finally, the parent (or canonical) configuration is used to carry out the numerical integration over the volume of the element. The elements can be iso-parametric, or, for simple cases, the mapping between the parent and the reference configuration can be an affine transformation.
@@ 2.3.2 ANCF 3443 Shell Element
Let $\mathbf{X}_{a}$ and $\mathbf{r}_{a}$ denote the nodal positions in the reference and current configurations, respectively, and let $\mathbf{d}_{a}=\mathbf{r}_{a}-\mathbf{X}_{a}$ be the displacement of node $a$. The isoparametric mappings for the reference position and displacement fields inside the element are
@@
The current position is $\mathbf{r}=\mathbf{X}+\mathbf{u}$, and the deformation gradient in the Total Lagrangian setting is obtained from
$\mathbf{F}(\xi,\eta,\zeta)=\frac{\partial\mathbf{r}}{\partial\mathbf{X}}=\frac{\partial\mathbf{r}}{\partial(\xi,\eta,\zeta)}\left(\frac{\partial\mathbf{X}}{\partial(\xi,\eta,\zeta)}\right)^{-1}.$ (10)
@@ 3 Augmented Lagrangian Formulation
$\quad+h\,\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v})^{T}\bigl(\bm{\lambda}+\rho\,\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})\bigr)\;,$
where $\mathbf{M}$ is a preassembled and constant mass matrix of the dynamic system, $\rho$ is the ALM penalty parameter, and $\mathbf{c}(\mathbf{q})=0$ denotes bilateral constraints with constraint Jacobian $\mathbf{C}_{q}=\partial\mathbf{c}/\partial\mathbf{q}$. Note that $\mathbf{c}(\mathbf{q})\in\mathbb{R}^{m}$ is a constraint-space quantity and is mapped to the generalized space through $\mathbf{C}_{q}(\cdot)^{T}$ in Eq. (17). We evaluate $\mathbf{C}_{q}^{T}(\cdot)$ via a sparse representation of $\mathbf{C}_{q}^{T}$, and enforce constraints with an augmented-Lagrangian outer loop that alternates an inner solve for $\mathbf{v}$ with multiplier updates until $\|\mathbf{c}\|_{2}$ meets a prescribed tolerance.
@@
Using $\nabla_{\mathbf{v}}\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})=h\,\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v})$ and the chain rule, one verifies that $\nabla_{\mathbf{v}}\Phi_{\rho}(\mathbf{v},\bm{\lambda})=\mathbf{g}(\mathbf{v},\bm{\lambda})$. Consequently, for fixed multipliers $\bm{\lambda}$, the inner solve $\mathbf{g}(\mathbf{v},\bm{\lambda})=\mathbf{0}$ is equivalent to enforcing stationarity of the augmented cost (18) with respect to $\mathbf{v}$, while the outer ALM loop updates $\bm{\lambda}$ to reduce the constraint violation $\|\mathbf{c}\|_{2}$.
@@ 6.2.1 Kinematics and Strain-Rate Measure in the Reference Configuration
In the Total Lagrangian setting, all kinematic quantities are expressed with respect to the reference configuration. Let $\mathbf{X}$ denote a material point in the reference configuration and $\mathbf{r}(\mathbf{X},t)$ its current position at time $t$. The deformation gradient is
$\mathbf{F}(\mathbf{X},t)=\frac{\partial\mathbf{r}}{\partial\mathbf{X}},\qquad J=\det\mathbf{F}.$ (55)
@@
$\dot{\mathbf{F}}=\frac{\partial\dot{\mathbf{r}}}{\partial\mathbf{X}},$ (57)
@@ 8 Kinematic Constraint Handling in TL-FEA Formulation
We summarize the representation and enforcement of bilateral kinematic constraints in the proposed TL-FEA multibody setting. We employ a compact set of scalar geometric primitives—dot-product 1 (DP1), dot-product 2 (DP2), distance (DIST), and coordinate-difference (CD)—as building blocks for intermediate constraints and engineering joints obtained through constraint composition. Practical applications may involve additional primitives and joint-specific constraints beyond those listed here; we therefore restrict attention to these four canonical forms to present the essential theory. The same point-evaluation and linearization operators (87)–(88) extend directly to richer constraint libraries. Each primitive is posed as a scalar holonomic condition at the position level, and its first variation provides the Jacobian contributions required by the velocity-level augmented Lagrangian residual in Eq. (17). To keep the presentation focused, we first define a unified constraint interface consistent with the step map $\mathbf{q}_{n+1}=\mathbf{q}_{n}+h\,\mathbf{v}$, then detail the four primitives in TL-FEA notation, and finally outline joint construction by composition together with enforcement within the ALM outer loop.
8.1 Constraint interface and TL-FEA evaluation operators
Let $\mathbf{q}\in\mathbb{R}^{n_{q}}$ denote the global vector of nodal unknowns assembled over all deformable bodies. A set of $m$ bilateral holonomic constraints is posed at the position level as
$\mathbf{c}(\mathbf{q},t)=\mathbf{0},\qquad\mathbf{c}:\mathbb{R}^{n_{q}}\times\mathbb{R}\rightarrow\mathbb{R}^{m},$ (84)
with Jacobian
$\mathbf{C}_{q}(\mathbf{q},t)=\frac{\partial\mathbf{c}}{\partial\mathbf{q}}\in\mathbb{R}^{m\times n_{q}}.$ (85)
In the velocity-level time stepping used to obtain Eq. (17), constraints are evaluated at $\mathbf{q}_{n+1}=\mathbf{q}_{n}+h\,\mathbf{v}$, so that the constraint residual and Jacobian entering the nonlinear solve are $\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})$ and $\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})$.
The constraint contribution appearing in Eq. (17) is written in augmented form as
$h\,\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})^{\mathsf{T}}\!\left(\bm{\lambda}+\rho\,\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})\right),$ (86)
where $\bm{\lambda}\in\mathbb{R}^{m}$ are Lagrange multipliers and $\rho>0$ is the ALM penalty parameter (scalar or componentwise). Constraints are scleronomic if $\mathbf{c}$ has no explicit time dependence and rheonomic otherwise (e.g., if driven by a prescribed function $f(t)$); in the present discrete setting, time dependence enters through evaluation of $\mathbf{c}(\cdot,t_{n+1})$.
Each primitive constraint is defined through the evaluation of points (and directions built from point differences) on one or more deformable bodies. For body $b$, element $E$, and material point $\mathcal{P}$ with reference coordinates $\mathbf{u}_{\mathcal{P}}$, the TL-FEA interpolation gives
$\mathbf{r}_{bE}(\mathcal{P};t)=\mathbf{N}_{E}^{b}(t)\,\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}}).$ (87)
For an element-level stacked unknown vector $\mathbf{e}_{E}^{b}=[\mathbf{e}_{1}^{\mathsf{T}},\ldots,\mathbf{e}_{n_{u}}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{3n_{u}}$, the corresponding linearization is
$\delta\mathbf{r}_{bE}(\mathcal{P};t)=\left(\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right)\delta\mathbf{e}_{E}^{b}.$ (88)
Directions used by the primitives are constructed from point differences; for two points $\mathcal{P}$ and $\mathcal{Q}$ we define $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$, where $\mathbf{r}_{\mathcal{P}}:=\mathbf{r}(\mathcal{P};t)$ and $\mathbf{r}_{\mathcal{Q}}:=\mathbf{r}(\mathcal{Q};t)$, so that $\delta\mathbf{a}=\delta\mathbf{r}_{\mathcal{Q}}-\delta\mathbf{r}_{\mathcal{P}}$ and Eq. (88) applies directly.
8.2 Primitive constraints in TL-FEA notation
The primitives DP1, DP2, DIST, and CD are scalar compositions of point evaluations (87) and point differences, so their Jacobians follow the same chain-rule pattern. For compactness, we state each primitive as a scalar constraint $c(\mathbf{q},t)=0$, provide its first variation $\delta c$, and summarize the corresponding nonzero element-level Jacobian blocks implied by Eq. (88).
8.2.1 Dot-product 1 (DP1)
DP1 prescribes the dot product between two directions, each defined by a pair of points. Let $\mathcal{P},\mathcal{Q}$ lie on body $b$ (elements $E,F$) and $\mathcal{R},\mathcal{T}$ lie on body $c$ (elements $G,H$). With the directions
$\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}},\qquad\mathbf{b}=\mathbf{r}_{\mathcal{T}}-\mathbf{r}_{\mathcal{R}},$ (89)
DP1 enforces
$c_{\mathrm{DP1}}(\mathbf{q},t)=\mathbf{a}^{\mathsf{T}}\mathbf{b}-f(t)=0,$ (90)
where $f(t)$ is prescribed (often constant; $f\equiv 0$ yields a perpendicularity condition). The first variation is
$\delta c_{\mathrm{DP1}}=\mathbf{b}^{\mathsf{T}}\delta\mathbf{a}+\mathbf{a}^{\mathsf{T}}\delta\mathbf{b},\qquad\delta\mathbf{a}=\delta\mathbf{r}_{\mathcal{Q}}-\delta\mathbf{r}_{\mathcal{P}},\quad\delta\mathbf{b}=\delta\mathbf{r}_{\mathcal{T}}-\delta\mathbf{r}_{\mathcal{R}}.$ (91)
Using (88) at each point, the only nonzero Jacobian blocks are those associated with the four hosting elements. For example, the contribution associated with point $\mathcal{P}$ has the form
$\frac{\partial c_{\mathrm{DP1}}}{\partial\mathbf{e}_{E}^{b}}=-\mathbf{b}^{\mathsf{T}}\left(\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),$ (92)
and the blocks associated with $\mathcal{Q},\mathcal{R},\mathcal{T}$ follow by replacing $(E,\mathcal{P},-\mathbf{b})$ with $(F,\mathcal{Q},+\mathbf{b})$, $(G,\mathcal{R},-\mathbf{a})$, and $(H,\mathcal{T},+\mathbf{a})$, respectively.
8.2.2 Dot-product 2 (DP2)
DP2 prescribes the dot product between a body-attached direction and a connector direction. Let $\mathcal{P},\mathcal{Q},\mathcal{R}$ lie on body $b$ (elements $E,F,G$) and $\mathcal{T}$ lie on body $c$ (element $H$). Define
$\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}},\qquad\mathbf{b}=\mathbf{r}_{\mathcal{T}}-\mathbf{r}_{\mathcal{R}},$ (93)
where $\mathbf{a}$ is a direction on body $b$ and $\mathbf{b}$ connects point $\mathcal{R}$ on body $b$ to point $\mathcal{T}$ on body $c$. DP2 enforces
$c_{\mathrm{DP2}}(\mathbf{q},t)=\mathbf{a}^{\mathsf{T}}\mathbf{b}-f(t)=0,$ (94)
with first variation
$\delta c_{\mathrm{DP2}}=\mathbf{b}^{\mathsf{T}}\delta\mathbf{a}+\mathbf{a}^{\mathsf{T}}\delta\mathbf{b},\qquad\delta\mathbf{a}=\delta\mathbf{r}_{\mathcal{Q}}-\delta\mathbf{r}_{\mathcal{P}},\quad\delta\mathbf{b}=\delta\mathbf{r}_{\mathcal{T}}-\delta\mathbf{r}_{\mathcal{R}}.$ (95)
The only nonzero Jacobian blocks are associated with the hosting elements of $\mathcal{P},\mathcal{Q},\mathcal{R}$, and $\mathcal{T}$, and are given by
$\frac{\partial c_{\mathrm{DP2}}}{\partial\mathbf{e}_{E}^{b}}$ $=-\mathbf{b}^{\mathsf{T}}\left(\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),$ $\frac{\partial c_{\mathrm{DP2}}}{\partial\mathbf{e}_{F}^{b}}$ $=+\mathbf{b}^{\mathsf{T}}\left(\mathbf{s}_{F}^{b}(\mathbf{u}_{\mathcal{Q}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),$
$\frac{\partial c_{\mathrm{DP2}}}{\partial\mathbf{e}_{G}^{b}}$ $=-\mathbf{a}^{\mathsf{T}}\left(\mathbf{s}_{G}^{b}(\mathbf{u}_{\mathcal{R}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),$ $\frac{\partial c_{\mathrm{DP2}}}{\partial\mathbf{e}_{H}^{c}}$ $=+\mathbf{a}^{\mathsf{T}}\left(\mathbf{s}_{H}^{c}(\mathbf{u}_{\mathcal{T}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right).$ (96)
8.2.3 Distance (DIST)
DIST prescribes the separation between two points. Let $\mathcal{P}$ lie on body $b$ (element $E$) and $\mathcal{Q}$ lie on body $c$ (element $F$), and define $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$. We employ the squared-distance form
$c_{\mathrm{DIST}}(\mathbf{q},t)=\frac{1}{2}\left(\mathbf{a}^{\mathsf{T}}\mathbf{a}-f(t)^{2}\right)=0,$ (97)
with $f(t)>0$. The first variation is
$\delta c_{\mathrm{DIST}}=\mathbf{a}^{\mathsf{T}}\delta\mathbf{a},\qquad\delta\mathbf{a}=\delta\mathbf{r}_{\mathcal{Q}}-\delta\mathbf{r}_{\mathcal{P}}.$ (98)
Therefore, the only nonzero Jacobian blocks are those for the hosting elements of $\mathcal{P}$ and $\mathcal{Q}$:
$\frac{\partial c_{\mathrm{DIST}}}{\partial\mathbf{e}_{E}^{b}}=-\mathbf{a}^{\mathsf{T}}\left(\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),\qquad\frac{\partial c_{\mathrm{DIST}}}{\partial\mathbf{e}_{F}^{c}}=+\mathbf{a}^{\mathsf{T}}\left(\mathbf{s}_{F}^{c}(\mathbf{u}_{\mathcal{Q}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right).$ (99)
8.2.4 Coordinate-difference (CD)
CD constrains one Cartesian component of a point difference. Let $\mathcal{P}$ lie on body $b$ (element $E$) and $\mathcal{Q}$ lie on body $c$ (element $F$), and define $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$. Given a constant unit vector $\mathbf{d}\in\{[1,0,0]^{\mathsf{T}},[0,1,0]^{\mathsf{T}},[0,0,1]^{\mathsf{T}}\}$ selecting the $x$, $y$, or $z$ component, CD enforces
$c_{\mathrm{CD}}(\mathbf{q},t)=\mathbf{d}^{\mathsf{T}}\mathbf{a}-f(t)=0.$ (100)
The first variation is
$\delta c_{\mathrm{CD}}=\mathbf{d}^{\mathsf{T}}\delta\mathbf{a},\qquad\delta\mathbf{a}=\delta\mathbf{r}_{\mathcal{Q}}-\delta\mathbf{r}_{\mathcal{P}},$ (101)
so that the only nonzero Jacobian blocks are
$\frac{\partial c_{\mathrm{CD}}}{\partial\mathbf{e}_{E}^{b}}=-\mathbf{d}^{\mathsf{T}}\left(\mathbf{s}_{E}^{b}(\mathbf{u}_{\mathcal{P}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right),\qquad\frac{\partial c_{\mathrm{CD}}}{\partial\mathbf{e}_{F}^{c}}=+\mathbf{d}^{\mathsf{T}}\left(\mathbf{s}_{F}^{c}(\mathbf{u}_{\mathcal{Q}})^{\mathsf{T}}\otimes\mathbf{I}_{3}\right).$ (102)
Vector coincidence constraints are obtained by stacking three CD equations with $\mathbf{d}=\mathbf{e}_{x},\mathbf{e}_{y},\mathbf{e}_{z}$.
Since Eq. (17) requires the transpose action $\mathbf{C}_{q}^{\mathsf{T}}(\cdot)$, we typically avoid forming $\mathbf{C}_{q}$ explicitly. Instead, each scalar constraint $c_{i}$ contributes a local scatter term of the form $\mathbf{C}_{q,i}^{\mathsf{T}}\eta_{i}$ with $\eta_{i}=\lambda_{i}+\rho\,c_{i}$, which mirrors standard FE assembly and integrates naturally with the element-level loops used for internal and external force accumulation.
8.3 Joint construction by composition and enforcement within ALM
Engineering joints are obtained by stacking primitive constraints into a vector constraint $\mathbf{c}=[c_{1},\ldots,c_{m}]^{\mathsf{T}}$. This approach is convenient in the TL-FEA setting because all defining points and directions are evaluated through (87); attachments and body-attached axes therefore evolve consistently with the underlying deformable discretization. A spherical joint is realized by enforcing coincidence of two attachment points, implemented by stacking three CD constraints to impose $\mathbf{r}_{\mathcal{P}}-\mathbf{r}_{\mathcal{R}}=\mathbf{0}$. A revolute joint is obtained by combining the spherical constraint with two independent orientation constraints (DP1/DP2-type) enforcing axis collinearity, thereby suppressing the two rotational degrees of freedom orthogonal to the hinge axis. A fixed joint is obtained by combining point coincidence with three independent orientation constraints that eliminate relative rotation. Other joints (e.g., cylindrical, planar, prismatic) follow by the same composition principle by combining coincidence or axis-alignment constraints with additional translational or rotational suppression constraints, and are omitted for brevity.
Constraint enforcement is integrated into the velocity-level residual solve in Eq. (17). At each time step, we evaluate $\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})$ and apply $\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v},t_{n+1})^{\mathsf{T}}$ in sparse form through the primitive-local scatter operations described above, while solving the inner nonlinear system for $\mathbf{v}$ with $\bm{\lambda}$ held fixed. The ALM outer loop then updates the multipliers until $\|\mathbf{c}\|_{2}$ meets a prescribed tolerance, using the standard update
$\bm{\lambda}^{(k+1)}=\bm{\lambda}^{(k)}+\rho\,\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v}^{(k)},t_{n+1}).$ (103)
At convergence, the multipliers $\bm{\lambda}$ are the discrete dual variables associated with enforcing $\mathbf{c}(\mathbf{q}_{n+1},t_{n+1})=\mathbf{0}$. The corresponding generalized constraint reaction is recovered as $\mathbf{C}_{q}(\mathbf{q}_{n+1},t_{n+1})^{\mathsf{T}}\bm{\lambda}$ (up to the adopted sign convention); in the present velocity-level discretization, $h\,\mathbf{C}_{q}^{\mathsf{T}}\bm{\lambda}$ represents the associated constraint impulse over the time step. For coincidence constraints expressed by three CD equations, the components of $\bm{\lambda}$ admit a direct Cartesian interpretation as the reaction intensity at the attachment.
9 Summary of Contributions
The current work makes the following contributions:
  • We develop a TL deformable-body dynamics formulation in which kinematics and constitutive response are written uniformly in terms of the deformation gradient $\mathbf{F}$ (and its consistent linearization), preserving the ANCF objective of avoiding rotation interpolation while adopting standard finite-strain continuum mechanics structure Bonet2008nonlinear ; Shabana1998 .
  • We reorganize Shabana’s ANCF kinematics by expressing the position field as $\mathbf{r}(\mathbf{X},t)=\mathbf{N}(t)\mathbf{s}(\mathbf{X})$ rather than $\mathbf{r}(\mathbf{X},t)=\mathbf{S}(\mathbf{X})\mathbf{e}(t)$, which shifts the placement of shape functions and yields the explicit factorization $\mathbf{F}(\mathbf{X},t)=\mathbf{N}(t)\mathbf{H}(\mathbf{X})$ with $\mathbf{H}=\partial\mathbf{s}/\partial\mathbf{X}$ shabanaYakoub2001Part1 ; Shabana2023Overview .
  • We derive internal forces and consistent tangent operators directly in deformation-gradient form, separating precomputable geometric operators from material response and enabling reuse across ANCF beam/shell elements and standard solid discretizations, as well as straightforward incorporation of additional finite-strain material models Gerstmayr2013 ; Otsuka2022RecentAdvances .
  • The resulting framework constitutes a unified formulation layer for systems of interacting deformable bodies, where element topology enters only through basis definition and reference parameterization, while the governing equations and linearizations follow the same $\mathbf{F}$-based template Gerstmayr2013 ; Shabana2023Overview .